OpenStudy (anonymous):

Trig Identities Prove: \[\sin^4x+\cos^4x=\sin^2x(\csc^2x-2\cos^2x)\]

ganeshie8 (ganeshie8):

write \(\sin^4x+\cos^4x\) as \((\sin^2x)^2 +(\cos^2x)^2\) and recall the identity \(a^2+b^2=(a+b)^2-2ab\)

ganeshie8 (ganeshie8):

\[\begin{align}\sin^4x+\cos^4x &=(\sin^2x)^2+(\cos^2x)^2\\~\\ &=(\sin^2x+\cos^2x)^2-2\sin^2x\cos^2x\\~\\ &=(1)^2-2\sin^2x\cos^2x\\~\\ &=1 - 2\sin^2x\cos^2x\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

factor out \(\sin^2x\)

OpenStudy (anonymous):

I got it now! Thank you so much!

ganeshie8 (ganeshie8):

yw